# State Frequency Estimation for Anomaly Detection

### v1

@@ VII-C SEQUENT’s Robustness
Tables V presents SEQUENT’s robustness against the aforementioned adversarial attacks on each selected dataset. The results are the average AUC scores across 10 independent runs. Based on the reported results, we see that SEQUENT achieved the best performance against two types of adversarial attacks (Padding Attack and Random Replacement Attack) compared to all other models. This can be explained by the nature of these two attacks; when an adversary tries to mimic the normal network behavior by padding or doing replacements on their flows, the traces will resemble the traces observed at training. However, this also changes the sequential behavior exhibited in the traces, which also results in changes in the state visit frequencies during the replay. SEQUENT detects this unusual increase in visit frequencies in the states and will raise an alarm. Existing methods trained on the raw feature value will not detect this deviation as these values are similar to what was observed during training. Probabilistic models such as Markov Chains and probabilistic state machines also suffer from this issue as the adversarial traces are similar to what was observed during training. The models will assign higher probabilities to these traces, thus resulting in poor detection performance.
@@
TABLE V: The performance (average AUC) of each model against the four adversarial attacks.
VIII Conclusion & Future Work
@@ Acknowledgments
The actual acknowledgment text is removed for authors’ anonymity and will be added after the review process. The current text is used as a placeholder so that we can take the required space into account later.

### v2

@@ 7.4. SEQUENT’s Robustness
Tables 7 presents SEQUENT’s robustness against the aforementioned adversarial attacks on each selected dataset. The results are the average AUC scores across 10 independent runs. Based on the reported results, we see that SEQUENT achieved the best performance against two types of adversarial attacks (Padding Attack and Random Replacement Attack) compared to all other models. This can be explained by the nature of these two attacks; when an adversary tries to mimic the normal network behavior by padding or doing replacements on their flows, the traces will resemble the traces observed at training. However, this also changes the sequential behavior exhibited in the traces, which also results in changes in the state visit frequencies during the replay. SEQUENT detects this unusual increase in visit frequencies in the states and will raise an alarm. Existing methods trained on the raw feature value will not detect this deviation as these values are similar to what was observed during training. Probabilistic models such as Markov Chains and probabilistic state machines also suffer from this issue as the adversarial traces are similar to what was observed during training. The models will assign higher probabilities to these traces, thus resulting in poor detection performance.
@@
Table 7. The performance (average AUC) of each model against the four adversarial attacks.
8. Conclusion & Future Work
@@ Acknowledgments
This work is funded under the Assurance and certification in secure Multi-party Open Software and Services (AssureMOSS) Project, (https://assuremoss.eu/en/), with the support of the European Commission and H2020 Program, under Grant Agreement No. 952647.